40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Sep 11 00:47:14 2014, Evan, DailyProgress, optic, South photothermal TF spt.pdfsouthPT.zipphotothermalTF.pdf
    Reply  Thu Sep 11 11:40:41 2014, Evan, DailyProgress, optic, South photothermal TF spt.pdfsouthPT.zipphotothermalTF.pdf
Message ID: 1512     Entry time: Thu Sep 11 11:40:41 2014     In reply to: 1511
Author: Evan 
Type: DailyProgress 
Category: optic 
Subject: South photothermal TF 

Quote:

I took a swept-sine measurement of the photothermal TF just as Tara and I did for the north cavity. To get a better measurement, I made some configuration changes:

  • I turned the power incident on the south cavity up to 8.5 mW by adjusting the post-laser HWP from 318° to 286°.
  • I placed an OD2.0 in front of the beat PD to prevent RF saturation.

Settings/values:

  • The beat was at 13.8 MHz.
  • The PLL Marconi was on 50 kHz FM deviation, and the SR560 gain was 100 V/V.
  • South transmission PD was 460(5) mV dc.
  • South transmission power (directly out of vacuum chamber) was 2.20(5) mW dc.

The results are attached. I'm not sure why there's a discrepancy around 200 Hz between the two traces. Below 100 Hz the measurement looks relatively clean.

The light rejected out of the post-EOAM PBS is only 2 mW (compared with 9 mW transmitted), which makes me suspicious that the post-EOAM QWP is not rotated properly, or else the input polarization into the EOAM is wrong. We should check this before redoing this measurement.

As with the north cavity, I find that an absorption of 6 ppm is needed make the measured curve lie on top of the theory curve.

For the time being, I have left the input power at 8 mW in case we want to take this again tomorrow. There's currently a dump upstream of the PMC to block the beam.

I didn't like the EOAM situation, so I rotated the post-EOAM QWP from 302° to 285°. With no voltage applied to the EOAM, this gives 6 mW of p and 6 mW of s. This may not be the true optimal setting, but the previous 2 mW / 9 mW situation seems too weird to be right. For commissioning the ISS I suspect we'll have to redo this EOAM setup to make sure the polarizations are behaving as we think.

The results are attached. I'm still seeing discrepancies at the points where the TFs are stitched together. Maybe it's because I'm using the SR785's auto source adjust feature.

 

Attachment 1: spt.pdf  65 kB  | Hide | Hide all
spt.pdf
Attachment 2: southPT.zip  14 kB
Attachment 3: photothermalTF.pdf  102 kB  | Hide | Hide all
photothermalTF.pdf
ELOG V3.1.3-